The girl has a speed of 7 m/s at the bottom of the ramp.

Calculate the kinetic energy of the girl at the bottom of the ramp.


1
Expert's answer
2021-03-02T18:07:11-0500
"KE=\dfrac{1}{2}mv^2,""KE=\dfrac{1}{2}\cdot50\ kg\cdot(7\ \dfrac{m}{s})^2=1225\ J."